| Post/Author/DateTime | Post |
|---|---|
| Fwib02-21-08, 11:05 AM | [edit] Just realised that you only get one actual attack with palm throw, so that makes it much more difficult to get near-endless attacks. :( Ah well, I'll get you next time, Gadget! Have a pile of kukris. Take Improved Critical, to shift the crit-rate up to 15-20 Take Lightning Mace (and its prereqs) (Complete Warrior) Take Chakram Ricochet (usually needs EWP:Chakram) (Champions of Ruin) Use Warblade's Weapon Aptitude to make those last two usable on the Kukri. (both are feats whic apply to only a single weapon, so RAW, they're valid) Get hold of the Palm Throw trick, for throwing two weapons per attack (Complete Warrior, Master Thrower PrC) (does need the DM to agree that the kukri is a 'little thrown weapon') Get Bloodstorm Blade 4, for the weapon-returning ricochet ability. (Tome of Battle) Be in Blood In The Water stance (also ToB) Palm-Throw a pair of kukris, and ricochet them onto adjacent foes from the targets. That's potentially four attacks using up only one of your full-attack. Each of those four attacks has a 30% chance to threaten, causing an extra attack from Lightning MaceKukri - total 120117% change for an extra attack, from one attack - probably unlimited attacks. Every actual crit you get (not just threats :( ) gives you +1/hit and damage from Blood in the Water. OK. 1) If this has been posted before, please link to it for me, Thanks. 2) Clearly this could do with some improvement, relying as it does on adjacent opponents - otherwise you just get lots of attacks, not unlimited - know a fix for this? 3) Your DM may beat you up, but I think it's RAW-legal. Surreal's list of extra attacks (http://forums.gleemax.com/showthread.php?t=627923) may help, but I'm too lazy to do more right now. |
| Aceon02-21-08, 12:02 PM | Answer here (http://forums.gleemax.com/showpost.php?p=15141586&postcount=8) |
| Fwib02-21-08, 03:51 PM | Answer here (http://forums.gleemax.com/showpost.php?p=15141586&postcount=8)"The designers didn't intend it that way. It's a DM call" Is there a single thread on this board where the designers did intend it that way? Everything in the game is a DM call. (Although, admittedly, my DM did break down, curl into a ball, and start babbling BanBanBanBanBan) :) |
| Aceon02-21-08, 06:20 PM | Maybe, but I bet he would stand tall and regain his sanity if he saw the answer from Cust Serv. |
| Aceon02-21-08, 06:29 PM | You want unlimited attacks? You can do it legit with Bloodstorm Blade. Attack everyone within the area of your thrown melee weapon once. Throw on far shot feat and distance weapon and maybe one of those ranger spells that get you no range penalty and you get exactly one attack at everything within 100 feet or so. In a pitched battee with an army of orcs, this could amount a whole lot of attacks. Presents a pretty cool visual too. One guy against an army. :D |
| Fwib02-21-08, 06:48 PM | You want unlimited attacks? You can do it legit with Bloodstorm Blade. Attack everyone within the area of your thrown melee weapon once. Throw on far shot feat and distance weapon and maybe one of those ranger spells that get you no range penalty and you get exactly one attack at everything within 100 feet or so. In a pitched battee with an army of orcs, this could amount a whole lot of attacks. Presents a pretty cool visual too. One guy against an army. :DOoh. Bag-o-rats! |
| altpersona02-21-08, 10:41 PM | for unlimited attacks... combine the 'attack everyone once' and a dragonwroght kobold / epic range feat so you have no range penalty... your range = line of sight.. requires spot of 20 ranks (17th lvl, might as well actually be epic..)... so that bites.. |
| Bauglir02-21-08, 11:35 PM | Each of those four attacks has a 30% chance to threaten, causing an extra attack from Lightning MaceKukri - total 120% change for an extra attack, from one attack - probably unlimited attacks. Probability does not work that way! Unless I'm misremembering horribly, it'd be 70%^4 chance NOT to critical, or 24.01%. So you'd still have a better than 3/4 chance of critting and getting another attack, but you'll fail eventually. |
| Fwib02-22-08, 11:38 AM | Probability does not work that way! Unless I'm misremembering horribly, it'd be 70%^4 chance NOT to critical, or 24.01%. So you'd still have a better than 3/4 chance of critting and getting another attack, but you'll fail eventually.Does work that way. On average, the number of extra attacks per palm throw, assuming that that attack bonus is high enough to hit on a 2 (which it will be, due to Blood in the Water and a pile of crits), will be 117% (3% loss due to natural 1s on the first attacks causing loss of the ricochet. The chance of not getting any threats at all on a full-attack of palm throws and ricochets with a reasonable attack bonus is very small indeed (less than 1%) Certainly, if you start a combat with no BitW bonus, and enemies with very high ACs compared to your Attack Bonus, or enemies with miss chances, or only one enemy, then this is much less effective. |
| AncientDragon02-22-08, 12:33 PM | This (http://forums.gleemax.com/showthread.php?p=14130294#post14130294) can lead to unlimited strenght and it means unlimited damage and attack bonus. Add Cleave, then attack a foe (oviusly with unlimidet strenght you will kill him) then thanks to clave you'll have another attack. And then another, and another... |
| BetaSquirrell02-22-08, 02:45 PM | What happens if odd things with Summoning spells come into play? Like if you enchant your weapon so that whenever it strikes it casts Summon Monster I? If I understand C.R. correctly you would be able to bounce it into the summoned creatures, creating more creatures... |
| Bauglir02-22-08, 07:47 PM | Wait, wait. The way I'm reading it, you have a 30% chance to critical on your first attack, right? I'm assuming you auto-confirm. If you do, it doesn't matter what happens later (you may get more attacks sure, but this eventually catches up with you). So you go to your next attack, which also has a 30% chance of criticalling. This is 30% of the original 70% left over from previous attempts, or 21%. And so on. When you sum the total of the first four iterations of this, you end up with 75.99% as your chance of successfully having made a crit in a given round. If you DID make a crit, you can just ignore it for terms of calculating probability; it gives you another attack, so it's like it never happened for chances. Probabilities on iterative attempts add multiplicatively, they don't just add. |
| vegetalss402-23-08, 03:43 AM | four attacks with 3/10 change to crit per strike so change to not crit is 7/10 so the change to not get a crit on all four is 7/10*7/10*7/10*7/10=2401/10000= 24.01% so the change to crit is 100%-24.01%=75.99 i just realised that that was the same as what Bauglir just said but i could not see that from the way he wrote it. |
| Fwib02-23-08, 07:58 PM | You only need to threaten with Lightning Mace. On average, you'll get 117% chance of an extra attack with each palm throw, assuming you hit on roll of a 2. Yes, the chance to not get any extra attacks at all is around 1/4, but on some throws, you'd get more than one bonus attack - which we assume you can take all of, with the kukris instant-returning to take the extra attacks after the ricochet, or quick-drawing more kukris to throw. |
| colm_l02-24-08, 09:08 AM | Does work that way. On average, the number of extra attacks per palm throw, assuming that that attack bonus is high enough to hit on a 2 (which it will be, due to Blood in the Water and a pile of crits), will be 117% (3% loss due to natural 1s on the first attacks causing loss of the ricochet. The chance of not getting any threats at all on a full-attack of palm throws and ricochets with a reasonable attack bonus is very small indeed (less than 1%) Fwib, Bauglir's probability analysis uses the correct methods. You can't have a probability higher than 1 (or 100%). The probability of one or more successful crits will be: 1-[(1-[crit range]/20)^{number of initial attack rolls}] This link (http://en.wikipedia.org/wiki/Binomial_distribution) should explain how to calculate your chance of successes. |
| Fwib02-24-08, 12:44 PM | Fwib, Bauglir's probability analysis uses the correct methods. You can't have a probability higher than 1 (or 100%). The probability of one or more successful crits will be: 1-[(1-[crit range]/20)^{number of initial attack rolls}] This link (http://en.wikipedia.org/wiki/Binomial_distribution) should explain how to calculate your chance of successes.Please read what it says, not what you think it says. |
| Fwib02-24-08, 01:04 PM | I just noticed that palm-throw is only one attack, not two, which cuts this down to size. :( |
| colm_l02-24-08, 02:08 PM | Please read what it says, not what you think it says. I do research using these distributions for a living & recently had to teach a tutorial on them. If it doesn't say what I think it says my students are screwed (& so am I). ;) http://en.wikipedia.org/wiki/Binomial_probability was in fact what I should have linked- simpler non technical explanation. The maths for it is as follows (for 4 attacks): p=0.3 1-p=0.7 general form: f(r)=(nCr)[p^r]*[(1-p)^(n-r)] 0 criticals (4C0)=1 [.3^0]=1 [.7^4]=0.2401 f(0)=0.2401 1 critical (4C1)=4 [.3^1]=0.3 [.7^3]=0.343 f(1)=0.4116 2 criticals (4C2)=6 [.3^2]=0.09 [.7^2]=0.49 f(2)=0.2646 3 criticals (4C2)=4 [.3^3]=0.027 [.7^1]=0.7 f(3)=0.0756 4 criticals (4C4)=1 [.3^4]=0.0081 [.7^0]=1 f(4)=0.0081 Sum(f(1):f(4))=0.7599 |
| Fwib02-24-08, 04:02 PM | I do research using these distributions for a living & recently had to teach a tutorial on them. If it doesn't say what I think it says my students are screwed (& so am I). ;) http://en.wikipedia.org/wiki/Binomial_probability was in fact what I should have linked- simpler non technical explanation. The maths for it is as follows (for 4 attacks): p=0.3 1-p=0.7 general form: f(r)=(nCr)[p^r]*[(1-p)^(n-r)] 0 criticals (4C0)=1 [.3^0]=1 [.7^4]=0.2401 f(0)=0.2401 1 critical (4C1)=4 [.3^1]=0.3 [.7^3]=0.343 f(1)=0.4116 2 criticals (4C2)=6 [.3^2]=0.09 [.7^2]=0.49 f(2)=0.2646 3 criticals (4C2)=4 [.3^3]=0.027 [.7^1]=0.7 f(3)=0.0756 4 criticals (4C4)=1 [.3^4]=0.0081 [.7^0]=1 f(4)=0.0081 Sum(f(1):f(4))=0.7599 When I said 'read what it says' I meant 'read what what I wrote says' If you have four attack rolls, each with a 30% chance to threaten, then you get an average of more than one threat. [edit] I looked at your numbers. You forgot to multiply the probability of getting a certain number of threats by the number of threats before you added them together to get the expected number of threats. |
| colm_l02-24-08, 05:02 PM | When I said 'read what it says' I meant 'read what what I wrote says' If you have four attack rolls, each with a 30% chance to threaten, then you get an average of more than one threat. Ah right, you meant the mean. That's grand but it doesn't work for iterative procedures like this one. The weighting isn't accurate due to lightning mace If you use the method I was suggesting, you can calculate your expected number of free attacks off it. It won't be infinite but it should be quite high. [edit] I looked at your numbers. You forgot to multiply the probability of getting a certain number of threats by the number of threats before you added them together to get the expected number of threats. Hmmm. I didn't forget - I didn't bother. If I was doing an iterative routine to find the total number of attacks I would have done that. I was showing the math for what bauglir had said earlier. |